How to find d max height reached above d horizontal in an inclined projection

Grade:11

1 Answers

Sachet
7 years ago
Applying v^2 =u^+2as in vertical direction between the point of projection and the topmost point we get 0^2 =u^2sin^thetha-2gH H=u^2sin^2thetha/2gHope it clears your doubt